OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (rc94):

You could do simultaneous equations. You know a horse has four legs and one head, and a human has 2 legs and one head. If we let X be the number of horses and Y be the number of humans we have two equations: X + Y = 74 4X + 2Y = 196 or similarly 2X + Y = 98 and solve

OpenStudy (anonymous):

so assumeing that HM = Human and HR = Horse HM + HR = 74 2HM + 4HR = 196 (2HM + 4HR) - (2 HM + 2HR) = 196 - 148 2HR = 48 HR = 24 HM + (24) = 74 HM = 74 - 24 HM = 50 So, the solution is 24 horses and 50 humans ?

OpenStudy (anonymous):

is that right?

OpenStudy (rc94):

It is

OpenStudy (anonymous):

yay!